Derek scores 100 and 60 on his first two exams.what must he score for a minimum average of 70.write an inequality an solve
Found 2 solutions by stanbon, fractalier:
Answer by stanbon(75887)

Derek scores 100 and 60 on his first two exams.what must he score for a minimum average of 70.write an inequality an solve
---
(100 + 60 + x)/3 >= 70
----
160+x >= 210
---
x >= 50
-----------
Cheers,
Stan H.
-----------

Answer by fractalier(6550)

Call his third test score, x. Then we write
(100 + 60 + x)/3 >= 70
(160 + x) / 3 >= 70
160 + x >= 210
x >= 50
He must score at least 50.